OpenStudy (anonymous):

3tan^2 x+5tan x+2=0 I need the exact value for x=? and the approximate value for x=? plessssssssssss

OpenStudy (dumbcow):

solve it like a quadratic equation u = tan x --> 3u^2 +5u +2 = 0

OpenStudy (anonymous):

I tryed it but I couldnt get any silution

OpenStudy (dumbcow):

were you able to factor it?

OpenStudy (anonymous):

no :(

OpenStudy (dumbcow):

ok do you know the quadratic formula?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i did it

OpenStudy (anonymous):

it is either -5+ squr1/6 or it is either -5- squr1/6

OpenStudy (anonymous):

is that right ?

OpenStudy (dumbcow):

haha nevermind i misread your answer

OpenStudy (anonymous):

yeah after u solve it it will come up to -5+ squr1/6 or it is either -5- squr1/6

OpenStudy (anonymous):

hahahahahaha

OpenStudy (dumbcow):

ok well sqrt(1) is just 1 so you get -4/6 = -2/3 -6/6 = -1

OpenStudy (anonymous):

so do u know what is the exact value for x and what is the approximate value ?

OpenStudy (dumbcow):

to find that we need to sub back in tan(x) for u \[\tan(x) = -2/3\] \[\tan(x) = -1\]

OpenStudy (anonymous):

I loveeeeeeeeeeeeeeeeeeee uuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuu

OpenStudy (dumbcow):

use calculator and use inverse tan

OpenStudy (dumbcow):

alright then :)